OpenStudy (anonymous):

Solve for y: -7/5y -4/5=1/3y-3

OpenStudy (anonymous):

\[\frac{- 7 }{ 5 }y \frac{ -4 }{ 5 }=\frac{ 1 }{ 3 }y-3\]